OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x + sin x

hartnn (hartnn):

cos 2x=2 cos^2 x-1

OpenStudy (anonymous):

cos 2x= co(x+x)= cos x cos x - sinx sin x

OpenStudy (anonymous):

i think it is 1 +3 sin^2x

OpenStudy (anonymous):

cos 2x= co(x+x)= cos x cos x - sinx sin x =1- 2 sin^2 x = 2 cos^2x -1

hartnn (hartnn):

there are 3 different formulas for cos 2x,use any to bring it in the form of sin x and cos x then there will be 3 equivalent answers.

OpenStudy (anonymous):

cos2x=2sinxcocsx 2sinxcosx+sinx=sinx(2cosx+1)

OpenStudy (anonymous):

so what would the anserw be?

OpenStudy (anonymous):

1 + 2 sin^2x + sin x

hartnn (hartnn):

if u have choices, list them else any of the 3 answers are correct 1-2sin^2 x+sin x

OpenStudy (anonymous):

A1 + 3 sin x B1 + 3 sin2x C1 - 2 sin^2x + sin x D1 + 2 sin^2x + sin x

hartnn (hartnn):

2cos^2x -1+sin x and cos^2 x-sin^2x +sin x

hartnn (hartnn):

C matches.

OpenStudy (anonymous):

THANKS SO MUCH

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

cos^2x-sin^2x+sinx=1-sin^2x-sin^2x+sinx=1-2sin^2x+sinx

OpenStudy (anonymous):

babydoll u can understant that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ya ok